OpenStudy (anonymous):

Write an equation of a line that passes through (9,-5) and (3,-5). Hint: To answer this question, first find the slope using these 2 points.

OpenStudy (anonymous):

Find slope. \[m = \frac{ y2-y1 }{ x2-x1 } \to \frac{ -5 - (-5) }{ 3 - 9 }\]

OpenStudy (anonymous):

\[\frac{ -5 + 5 }{ 3 - 9 }\]

OpenStudy (anonymous):

ok :) hold on :)

OpenStudy (anonymous):

wait i got 0

OpenStudy (anonymous):

So you have a slope of 0. You could have also known this by checking the y-values in both points, which are the same.

OpenStudy (anonymous):

so the equation will look likw what

OpenStudy (anonymous):

The equation of a horizontal line is y =

OpenStudy (anonymous):

y = 0

OpenStudy (anonymous):

No, 0 is the slope. What is the y-value?

hartnn (hartnn):

lol! rebecca left without getting right answer, she is already on next question.....

OpenStudy (anonymous):

Yep! Let's see if she'll come back @rebecca1233 (:

hartnn (hartnn):

nopes, she won't come back i guess....leave it.